Circular No. 7055

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998eu AND 1998ev IN ANONYMOUS GALAXIES

A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise Observatory Optical Transients Search (see IAUC 6917) their discovery of two apparent supernovae on unfiltered CCD images taken with the Wise Observatory 1-m telescope on Nov. 14. SN 1998eu (R = 19.7) is located at α = 0^{h}59^{m}58^{s}.66, δ = +14^{o}18′00′′.4 (equinox 2000.0), which is in the field of the galaxy cluster Abell 125 (z = 0.188). SN 1998ev (R = 21.3) is located at α = 1^{h}04^{m}12^{s}.68, δ = +25^{o}00′59′′.1, which is in the cluster Abell 136 (z = 0.157). The host galaxies of both supernovae are only marginally detected, so no offsets can be given at this time. Both supernovae were invisible in images obtained in July (limiting mag R ∼ 22.5), but were confirmed on CCD images taken by J. Dan and P. Ibbetson on Nov. 17 with the same telescope. Finding charts can be ob-

tained via ftp://wise3.tau.ac.il/pub/avishay/sn1.fits (SN 1998eu)
and /sn2.fits (SN 1998ev), or via e-mail to avishay@wise.tau.ac.il.

PKS 2005–489 G. Tagliaferri, G. Ghisellini, and L. Maraschi, Osservatorio Astro-

nomico di Brera, Milan; and P. Giommi, BeppoSAX, Rome, on behalf of the BeppoSAX ToO collaboration on blazars, report: “This BL Lac object was observed with the BeppoSAX Narrow Field Instruments from Nov. 1.318 to 2.453 UT. The source was bright and well visible in the Phoswich Detector System up to 100 keV. The preliminary analysis indicates average energy fluxes of 5.4 and 1.7 × 10^{−10} erg cm^{−2} s^{−1} in the bands 0.1–10 and 10–100 keV, respectively. A broken power-law well describes the data, with photon indices 1.95 ± 0.03 and 2.21 ± 0.02 below and above 2 keV, respectively. This hard and bright x-ray flux suggests a high flux in the GeV-TeV band, with a maximum foreseen at ∼ 0.1 TeV on the basis of a simple synchrotron self-Compton model. Observations at other wavelengths are encouraged, in particular at the highest energies.”

COMET 21P/GIACOBINI-ZINNER Visual m_{1} and coma-diameter estimates: Oct. 25.74 UT, 8.9, 4′.4 (M.

Lehký, Hradec Králové, Czech Rep., 25×100 binoculars); Nov. 6.02, 8.5, 5′ (J. G. de S. Aguiar, 11×80 binoculars, Campinas, Brazil); 17.37, 8.7, 5′ (S. T. Rae, Te Awamutu, New Zealand, 10×50 binoculars); 20.08, 9.0, 2′.5 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector).

1998 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green